T ITAN Chart Text Shape Media Comment Insert Table 24. Which activity BEST describes how GDP is calculated? a. summing the ma need 24-27 answered plz
Answer #1

24. A
(Summation of market value of all final goods is GDP.)

25. B
(Growth rate = [(1400-1300)/1300]*100 = (100/13) = 7.69%)

26. B
(Only statement 2 is correct. Economic growth is growth is real GDP.)

27. A
(Importance of competitive and open markets is demonstrated.)